Find k for the probability P(-0.256<T<k)=0.59 , where v=27
Answer #1

Solution,

Using t table,  

v = 27

P( -0.256 < T < k) = 0.59

= P( T < k) - P( T < -0.256) = 0.59

=  P( T < k) - 0.3999 = 0.59

=  P( T < k) = 0.59 + 0.3999

= P( T < k) = 0.9899

= P( T < 2.468) = 0.9899

k = 2.468